OpenStudy (anonymous):

find a quadratic equation with roots (4+i) and (4-i) this is what i did and im stuck (x-(4+i)(x-(4-i) x^2-x(4-i)-x(4+i)-(4+i)(4-i)

OpenStudy (anonymous):

x^2-8x+7

OpenStudy (anonymous):

17*

OpenStudy (anonymous):

(x-(4+i) (x-(4-i) => (x-4-i)(x-4+i) So let (x-4) = a, then (a-i)(a+i) = a^2 - i^2 we know that i = root of -1, hence i^2 = -1 and -(-1) = 1 hence the equation becomes a^2 + 1 => (x-4)^2 + 1 = x^2-8x+16+1 = x^2-8x+17. Hope it helped. :)

OpenStudy (anonymous):

it really realli did help me understand better i got tht answer before bt my steps kept bothering me thank you so so so so much !!